# Survival Analysis with Adversarial Regularization

### v4

@@ VI-D Results
We analyze how adversarial perturbations impact the characteristics of the population survival curves.
@@ VI-D1 Comparison to SOTA
We show the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and -worst-case perturbations in Tables III and II, respectively. We empirically show that SAWAR through CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various -perturbation magnitudes. Thus, SAWAR has better generalization performance, better calibration, and better adversarial robustness by up to 150% compared to baseline adversarial training methods and SOTA NN-based SA models. More detailed results, along with the exact metric values for each of the worst-case and the FGSM adversarial attacks for each dataset are available in the code repository linked in the abstract.
VI-D2 Failure Rate Distribution
@@
Each instance in the dataset has its own survival curve . All instance’s survival curves are visualized as the credible interval, where is sampled from the dataset distribution . Thus, let us assume that the input covariates are random variables . Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates . Moreover, we treat as a stochastic process :
@@
We compute the statistical quantities of such as expectation, lower 5th quantile , and upper 95th quantile .
@@
Fig. 2 visualizes the credible interval and mean of the survival function stochastic process in Eq. 23. The 90% credible interval of the stochastic process changes with respect to the strength of the adversarial regularization method (Fig. 2). We observe the trend that the credible interval slightly narrows down as the strength of the adversarial regularization method increases (Fig. 2). However, this trend has an exception with AAE-Cox. We hypothesize that since AAE-Cox is not truly an adversarial robust method (i.e., no perturbations to the input covariates during training), and by introducing more parameters via an encoder of a GAN, there are more model parameters to attack. The variance of the failure rate decreasing leads to a more robust model against adversarial perturbations. For example, when , then there is a mode collapse such that . This is the case for point-estimates of the parameter derived via maximum likelihood estimation. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 19, we introduce a trade-off between individualized survival curves with vulnerability to adversarial perturbations and a population survival curve that is completely resilient to adversarial perturbations. In doing so, while exhibiting significantly higher adversarial robustness, SAWAR’s credible interval remains relatively unchanged compared to the DRAFT method.
@@ VI-D3 Population Survival Curve
The -perturbation magnitude determines the severity of the worst-case survival curve. We observe the trend that the stronger the adversarial regularization method, the “closer” the -worst-case population curve becomes to the unperturbed population survival curve for a given (Fig. 3). Moreover, we use the Kaplan Meier Curve (KMC), a commonly used frequentist approach for non-parametrically estimating the population survival curve using samples , to visually evaluate model calibration. The close alignment between the population survival curve and the KMC shown in Fig. 3 demonstrates the strong calibration of SAWAR model [32].
@@ VII Conclusions
This paper introduces a robust training objective based on CROWN-IBP regularization to enhance adversarial robustness in fully-parametric NN-based SA models, improving both model calibration and generalization. We evaluated the SAWAR method on 10 time-to-event medical datasets, comparing it with standard adversarial regularization methods like Gaussian noise, FGSM, and PGD, as well as SOTA models such as DRAFT and AAE-Cox.
Our empirical results demonstrate that SAWAR consistently improves SA performance across various metrics, including CI, IBS, and NegLL, for all -perturbation magnitudes from 0 to 1. Compared to traditional adversarial regularization techniques, SAWAR shows superior resilience to input covariate perturbations, with relative improvements in CI and IBS ranging from 1
In summary, SAWAR significantly enhances predictive performance, adversarial robustness, and calibration, effectively mitigating data uncertainty in SA and improving generalization on unseen datasets.
@@
Appendix A Appendix

### v5

@@ VI-D Results
We analyze how input covariate perturbations impact the characteristics of the population survival curves, and SA metrics CI,IBS, and NegLL.
@@ VI-D1 Comparison to SOTA
We show the average ranking of each adversarial training method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and -worst-case perturbations in Tables III and II, respectively. We empirically show SAWAR leads to increased performance, consistently ranking the best across various -perturbation magnitudes. Thus, SAWAR has better generalization performance, better calibration, and better adversarial robustness by up to 150% compared to baseline adversarial training methods and SOTA NN-based SA models.
In order to confirm that SAWAR showed statistically significant improved performance, we performed a Friedman hypothesis test () for each metric (CI, IBS, and negative log-likelihood), using adversarial training methods as “treatments” and combinations of dataset, perturbation magnitude, and perturbation method as “blocks”. This test, suited for repeated measures, assessed whether our adversarial training method yielded more robust, well-fitted, and calibrated models across perturbation attacks compared to other adversarial training methods. All Friedman hypothesis tests had statistically signifcant results, which prompted post-hoc Conover-Iman test for pairwise comparisons amongst adversarial training methods, with p-values adjusted via the Benjamini-Hochberg procedure (critical difference diagrams shown in Fig. 2).
More detailed results, along with the exact metric values for each of the worst-case and the FGSM adversarial attacks for each dataset are available in Appendices D, E, F, A, B and C.
VI-D2 Survival Curve Distribution
@@
Each instance in the dataset has its own survival curve . Thus, let us assume that the input covariates are random variables . Then the hazard becomes a random variable that occurs from the stochasticity of the input covariates . Moreover, we treat as a stochastic process :
@@
Then, instance’s survival curves are visualized as the credible interval, where is sampled from the dataset distribution . We compute the statistical quantities of such as expectation, lower 5th quantile , and upper 95th quantile .
@@
Fig. 3 visualizes the credible interval and mean of the survival function stochastic process in Eq. 23. The 90% credible interval of the stochastic process changes with respect to the strength of the adversarial training method (Fig. 3). We observe the trend that the credible interval slightly narrows down as the strength (regularization) of the adversarial training method increases (Fig. 3). However, this trend has an exception with AAE-Cox. We hypothesize that since AAE-Cox is not truly an adversarial robust method (i.e., no perturbations to the input covariates during training), and by introducing more parameters via an encoder of a GAN, there are more model parameters to attack. The variance of the hazard decreasing leads to a more robust model against input covariate perturbations. For example, when , then there is a mode collapse such that . This is the case for point-estimates of the parameter derived via maximum likelihood estimation of a Poisson process without input covariates. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 19, we introduce a trade-off between individualized survival curves with vulnerability to input covariate perturbations and a population survival curve that is completely resilient to input covariate perturbations. In doing so, while exhibiting significantly higher adversarial robustness, SAWAR’s credible interval remains relatively unchanged compared to the DRAFT method.
@@ VI-D3 Population Survival Curve
The -perturbation magnitude determines the severity of the worst-case survival curve. We observe the trend that the stronger (regularization) the adversarial training method, the “closer” the -worst-case population curve becomes to the unperturbed population survival curve for a given (Fig. 4). Moreover, we use the Kaplan Meier Curve (KMC), a commonly used frequentist approach for non-parametrically estimating the population survival curve using samples , to visually evaluate model calibration. The close alignment between the population survival curve and the KMC shown in Fig. 4 demonstrates the strong calibration of SAWAR model [32].
Overall, robust and well-calibrated SA models are essential for making reliable clinical decisions, including evaluating treatment efficacy, analyzing dose-response relationships, determining optimal treatment timing, performing subgroup analysis, and assessing comparative effectiveness. In contrast, SA models prone to overfitting or sensitive to small perturbations in input covariates can yield unreliable or even harmful recommendations, potentially leading to fatal outcomes.
@@ VII Conclusions
This paper introduces SAWAR, a training objective that approximates a Min–Max optimization with CROWN-IBP to improve adversarial robustness in fully‑parametric NN-based SA models. We evaluated SAWAR on 10 benchmark medical time‑to‑event datasets and compared it to standard adversarial training methods (Gaussian noise, FGSM, PGD) and SOTA models (DRAFT, AAE-Cox). Our results show that SAWAR consistently improves performance across CI, IBS, and NegLL for various input covariate perturbation magnitudes . Compared with traditional adversarial training method, SAWAR offers superior resilience to input covariate perturbations, with relative gains in CI and IBS ranging from 1% to 150%. Overall, SAWAR enhances predictive accuracy, robustness, and calibration, effectively mitigating input covariate uncertainty and improving generalization to unseen datasets.
@@
Appendix A Concordance Index Worst-Case Perturbation
@@
Appendix B Integrated Brier Score Worst-Case Perturbation
@@
Appendix C Negative Log Likelihood Worst-Case Perturbation
@@ Appendix D Concordance Index FGSM Perturbation
Appendix E Integrated Brier Score FGSM Perturbation
@@ Appendix F Negative Log Likelihood FGSM Perturbation
Appendix G Friedman Hypothesis Test
Instead of relying on confidence intervals—which do not definitively determine whether one adversarial training method outperforms another—we conducted a Friedman hypothesis test. While it is common in the machine learning and AI research community to repeatedly perform pairwise hypothesis tests based on confidence intervals (e.g., repeatedly for each dataset comparing two adversarial methods performance), this approach has significant drawbacks. It increases the risk of a high false discovery rate, potentially leading to incorrect rejection of the null hypothesis, and often relies on the normality assumption, which may not hold in practice. Therefore, we conducted a Friedman hypothesis test at a significance level of 0.05 for each metric (CI, IBS, and negative log-likelihood), treating the adversarial training methods as the “treatments” and the combinations of datasets, perturbation strengths, and perturbation methods as the “blocks”. This test was chosen because it is well-suited for repeated measurements and allowed us to assess whether our training method produces a more robust, better-fitted, and more calibrated model across various perturbations (perturbation method and perturbation strength). Since the Friedman test revealed statistically significant differences for each metric, we conducted a post-hoc Conover-Iman test (a rank-based approach) to analyze pairwise differences between groups. To control the false discovery rate, we applied the Benjamini-Hochberg procedure for p-value adjustment. Our findings show that our method, SAWAR, is indeed statistical significant improvement in performance (see critical difference diagram in Fig. 5).